OpenStudy (anonymous):

what is the standard form of y=-3x/(4+5/4) I'll give a medal :D

OpenStudy (crashonce):

y=-12x / 9 therefore 9y=-12x therefore it is 12x+9y = 0 i think if i interpert the question correctly

OpenStudy (anonymous):

y=-3X/[4+5/4] y=3x*4/5*4 y=12x/20 y=3x/5

OpenStudy (anonymous):

3x/5 is the answer